OpenStudy (anonymous):

Which of the following is a solution of x2 - 10x = -36?

OpenStudy (solomonzelman):

add 36 to both sides, and factor.

OpenStudy (anonymous):

Set it equal to 0 (as a quadratic), factor, and see what your zeros/solutions are.

OpenStudy (anonymous):

so it would be x2-10x=36+0?

OpenStudy (anonymous):

\[x ^{2}-10x = -36\rightarrow x ^{2}-10x+36=0\]

OpenStudy (anonymous):

Factor: \[x ^{2}-10x+36\]

OpenStudy (anonymous):

6 and 4

OpenStudy (anonymous):

but those arent the answer choices

OpenStudy (anonymous):

they have to deal with i's

OpenStudy (anonymous):

Which of the following is a solution of x2 - 10x = -36? negative 5 minus i sqare root of 61 5 plus i square root of 61 negative 5 minus i square root of 11 5 plus i square root of 11

OpenStudy (anonymous):

No wait

OpenStudy (anonymous):

^ are the answer choices

OpenStudy (anonymous):

Yes, because x^2-10x+36 has no real roots

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

So, we must use the quadratic formula.

OpenStudy (anonymous):

Expect the discriminant to be (-)

OpenStudy (anonymous):

oh what joy im horrible at the quadratic formula

OpenStudy (anonymous):

Well in the form x^2-10x+36 We know a, b, and c You know the QF right? I'm not perfect with it either hehe

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

Would you like me to check your work as you go through it?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Alrighty =)

OpenStudy (anonymous):

Here, I'll set it up for you.